OpenStudy (anonymous):

Evaluate the expression for x = 3 and y = 6. 5(x2 + y) + y2

OpenStudy (alexwee123):

just plug in the values o.0 ...

OpenStudy (anonymous):

how

OpenStudy (alexwee123):

put in 3 for x and 6 for y

OpenStudy (anonymous):

77

OpenStudy (anonymous):

show me the steps lol

OpenStudy (anonymous):

5((3)2 + (6)) + (6)2 6 + 6 + 12 5 * 12 = 65 6 *2 = 12 65 + 12 = 77

OpenStudy (alexwee123):

@b.l.m is it X^2 or x*2?

OpenStudy (anonymous):

77 is the answer

OpenStudy (alexwee123):

well it can be also 111 if the x and y values have an exponents of 2

OpenStudy (anonymous):

Umm no, it's 77.

OpenStudy (alexwee123):

5(3^2 + 6) +6^2 5(15) +36 75+36 111

OpenStudy (anonymous):

Umm, alexwee is correct if that question reads as 5*(x-squared + y) + y-squared

OpenStudy (anonymous):

Well if it was x^2 then he would of put it.

OpenStudy (alexwee123):

well isn't it kind of weird that @b.l.m put the 2 on the right instead on the left of the variable?

OpenStudy (alexwee123):

so most likely it is x^2 and y^2 ....

OpenStudy (anonymous):

Umm well a letter always comes before a # unless you have a letter as an exponent. So sorry I'm right.

OpenStudy (anonymous):

But if that question reads as 5*(x*2 + y) + y*2, then the answer = 72

OpenStudy (anonymous):

No sorry its 77.

OpenStudy (anonymous):

Yeah 72.

OpenStudy (alexwee123):

and also if it wasn't squared then your answer is wrong anyways...

OpenStudy (alexwee123):

72 ya...

OpenStudy (anonymous):

A. 72 B. 87 C. 96 D. 111

OpenStudy (anonymous):

A

OpenStudy (alexwee123):

@b.l.m is it x^2 or x times 2??

OpenStudy (alexwee123):

because it can be either A or D....

OpenStudy (anonymous):

^2

OpenStudy (anonymous):

Then D.

OpenStudy (alexwee123):

OK then its D

OpenStudy (anonymous):

yep D

OpenStudy (anonymous):

next time b.l.m, type out the question better!

OpenStudy (anonymous):

okay thanks everyone